OpenStudy (anonymous):

Divide. –36 ÷ (–4) A. –40 B. –32 C. –9 D. 9

OpenStudy (anonymous):

It's nine

OpenStudy (anonymous):

\[\frac{-36}{-4}=\frac{(-1)\times36}{(-1)\times4}=\frac{(-1)}{(-1)}\times\frac{36}{4}=1\times\frac{9}{1}=9\]

OpenStudy (anonymous):

When multiplying or dividing by negative numbers, the answer always comes out to be the exact same as if you were multiplying or dividing the positive numbers. So -36 / -4 = 9. :) Hope this makes sense :P

OpenStudy (anonymous):

yeah, thank you, this helped me a lot :)

OpenStudy (anonymous):

Your welcome! :)

OpenStudy (anonymous):

No prob :)